The line through the point (1,0,6) and perpendicular to the plane "x+3y+z=5" has vector equation "\mathbf{r}=( u,v,w )=( a,b,c )+ t (d,e,f )" where t is a parameter.


Making use of the parameter t, find the values of: u; v and w



1
Expert's answer
2021-10-18T06:34:00-0400
"r=(1,0,6)+t(1,3,1)=(u,v,w)"

"u=1+t, v=3t, w=6+t"